HUBBLE SPACE TELESCOPE DAILY REPORT #2166 PERIOD COVERED: 0000Z (UTC) 07/15/98 - 0000Z (UTC) 07/16/98 Daily Status Report as of 197/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/1/2/3 7875 (Searching for Old Stellar Populations in Candidate Proto-Spheroidals) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations searching for old stellar populations in candidate proto-spheroid SA57-10601. We have compelling evidence from WF/PC-2 images that these galaxies have optical colors gradients. The near-infrared is necessary, however, to reveal color gradients large enough to separate young and old stellar populations. Only NICMOS imaging provides enough spatial resolution and depth to resolve these gradients. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 7910 (STIS Extragalactic Parallel Survey) The Space Telescope Imaging Spectrograph (CCD) was used to make extragalactic slitless spectrographic observations as part of the parallel high galactic latitude survey to search for emission line galaxies. Data was taken in parallel, and are not subject to proprietary time. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD/MA1 8016 (Effect of MAMA Charge Offsetting on Sensitivity and Dispersion Accuracy - Repeat) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of GD153. In order to conserve charge on the STIS MAMA microchannel plates and maintain even detector quantum efficiency, the STIS MAMA spectroscopic modes are differentially offset each month along and across the detector. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 7385 (Low-Mass Halo Binaries) The WF/PC-2 was used to make observations of the low mass, halo, binary, {Population II}, M-type, sub-dwarf, star LHS536. This star is too faint to be resolved through ground-based speckle interferometry, thus HST represents the only means of probing separations. The observations were completed as planned, and no anomalies were reported. 1.10 Completed NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.13 Completed STIS/CCD 7566 (Demographics of Nuclear Black Holes) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of normal galaxy NGC-7332 to try and detect a super massive black hole. A dark and wavelength calibration images were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Three Sets of NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. As described in HSTAR 6587, the STIS MAMA-2 MCE reset while in low voltage at 196/191419Z. Per ROP NS-12, the STIS flight software error counter was reset at 196/2004Z. MAMA-2 will be recovered via the SMS at the next time the high voltage is turned on, 197/171932Z. The Battery 1 Capacity Check Test continues. Discharge was terminated at 196/1435Z by the Special Hardware Limit Test. Recharge started at 196/1515Z, a normal orbit day entry event. Flight software will be configured back to a six-battery system approximately two days from the beginning of battery recharge. Routine reconditioning of ESTR-3 was performed at 197/0535Z per ROP RD-7. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH